OpenStudy (anonymous):

How many times does the graph of y = 1 intersect the graph of y = x2 + 1

OpenStudy (anonymous):

once at 0.

OpenStudy (anonymous):

let 1 = x^2 +1 x^2 = 0 x = 0

OpenStudy (anonymous):

y-1=x^2 it is an upward parabola. 

OpenStudy (solomonzelman):

https://www.desmos.com/calculator/raty7riwcs